# arXiv:2503.02268v1[cs.AI] 04 Mar 2025

### v1

@@ Abstract
Wenjia Jiang1,2, Yangyang Zhuang1, Chenxi Song1, Xu Yang3, Chi Zhang1, 1Westlake University, 2Henan University, 3Southeast University {jiangwenjia, chizhang}@westlake.edu.cn, https://appagentx.github.io/
@@ 2 Related Works
LLMs as GUI Agents. Beyond general agents, LLMs have also enhanced GUI automation, surpassing traditional script-based methods in flexibility and adaptability. WebVoyagerHe et al. (2024), AppAgentZhang et al. (2023), and MobileAgentWang et al. (2024b) leverage multimodal perception for interactive interfaces, while UFOZhang et al. (2024a), AutoGLMLiu et al. (2024), and MMAC-CopilotSong et al. (2024) improve cross-platform adaptability and multi-agent collaboration. To refine UI understanding, OmniParserLu et al. (2024) and Ferret-UIYou et al. (2024) enhance element recognition, while TinyClickPawlowski et al. (2024) and CogAgentHong et al. (2023) improve interaction precision and vision-based task execution. Additionally, WebGUMFuruta et al. (2024), MobileVLMSChu et al. (2024), and DigiRLBai et al. (2024) optimize performance in dynamic web and mobile environments. However, most existing agents rely on static training data and lack continuous adaptation. To address this limitation, we propose AppAgentX, integrating task trajectory memory to enhance efficiency and adaptability in GUI automation.
@@ 4.2 Evolutionary Mechanism
[…]
@@
4.3 Action Execution
Following the evolution of the action space, the agent can now select high-level actions from the expanded action space to efficiently execute tasks. The execution process begins after the agent captures a screenshot of the current page. At this stage, the system then matches the parsed elements on the page to the stored element nodes in memory, by comparing their visual embeddings. Next, we check whether these identified element nodes are associated with any shortcut nodes. If an association between the element nodes and shortcut nodes exists, the system leverages the LLM to determine whether the corresponding high-level actions can be executed. This decision is made by inspecting the description of the shortcut node in conjunction with the current task context. If the conditions for executing the high-level action are met, the LLM generates an action execution template. This template includes the sequence of low-level actions to be executed by the shortcut node, along with the corresponding function arguments necessary for each action. In cases where multiple elements on the page are associated with shortcut nodes, the system prioritizes executing the actions based on the order of matching elements. This order is determined by their execution sequence, ensuring that the actions are performed in a logical and efficient manner. If the conditions for executing a high-level action are not met, the agent will default to selecting actions from the basic action space $\mathcal{A}$. This ensures that the agent can still operate effectively, even when high-level abstractions cannot be applied to the current task.
@@ 5 Experiments
[…]
GPT-4o (Baseline) None Basic 10.8 26 6.72 16.9%
Element Basic 9.3 24 8.46 69.7%
AppAgentX Chain Basic 9.1 23 9.26 70.8%
Chain Basic+Evolve 5.7 16 4.94 71.4%
Table 1: Analysis of Different Components in AppAgentX. This table compares the performance differences resulting from the different designs with the baseline. Both our memory design and evolution mechanism can improve success rate and efficiency.
@@ 5.1 Experimental Setup
Mobile-Bench Deng et al. (2024): This benchmark offers a larger and more comprehensive set of task records, spanning multiple applications across 832 data entries. It provides a more extensive evaluation of the agent’s performance, facilitating comparisons with a broader range of existing methods in mobile task management.
Implementation details. The implementation of our framework leverages several key platforms. For the foundational LLM, we selected GPT-4o OpenAI (2024) as the default model unless otherwise stated. The LangGraph framework LangChain (2024) is used as the agent platform, providing essential features for LLM input-output parsing and process control. To implement the memory mechanism, we integrated Neo4j Neo4j (2024) for graph-based storage and retrieval and Pinecone Pinecone (2024) for vector search. For feature matching, we employed cosine similarity with embeddings derived from ResNet-50 He et al. (2015), which enables effective task representation and retrieval. All experiments were conducted using the Android Debug Bridge (ADB), enabling on-device evaluations for mobile applications.
[…]
DroidTask 158 AppAgent 106.24 11.5 46.3%
AppAgentX 56.29 5.1 88.2%
MobileBench (SAST) 332 AppAgent 150.24 7.6 72.6%
AppAgentX 42.38 5.4 86.1%
Table 2: Comparison with AppAgent on Large Benchmarks. This table evaluates the efficiency and accuracy of different frameworks on benchmarks containing a large number of tasks. In the MobileBench, SAST (Single-App-Single-Task) refers to a real dataset containing only a single task description
LLM Previous Sota AppAgentX
Gemini-1.5-Pro 25.6 12.4
Claude-3.5 38.4 11.4
GPT-4o 43.5 17.5
@@ 5.2 Experimental Analysis
Comparison with baselines. A comparison of our experimental results with model variants is shown in Table 1. We report the results on the AppAgent benchmark. We begin with a baseline model that does not incorporate any memory mechanism and progressively introduce our proposed enhancements. The first model variant integrates a module that records only information about elements, providing a basic form of memory. Subsequently, we introduce a more advanced memory design that maintains a structured memory chain, enabling the model to retain a more comprehensive representation of past interactions. Finally, we augment the system with an evolutionary mechanism that expands the action space to include high-level actions, further optimizing task execution.
@@
Task difficulties analysis. To further investigate the performance of AppAgentX across tasks of varying complexity, we categorize the tasks based on their ground-truth operation steps, annotated through expert demonstrations. Tasks with up to 5 steps are considered short, tasks with 6 to 10 steps are classified as medium-length, and tasks requiring more than 10 steps are categorized as long tasks. As shown in Figure 4, the violin plot provides an intuitive visualization of how task complexity affects performance. Notably, our method demonstrates a clear advantage in terms of task execution time as the task complexity increases. To account for random fluctuations, we conducted one-tailed paired t-tests Ross and Willson (2017) on each data group to verify that AppAgentX significantly outperforms the baseline in efficiency, with all p-values falling below 0.05. Our model consistently outperforms the baseline across all task lengths, resulting in reduced task completion times. This finding reinforces the robustness and efficiency of our approach, confirming its stability under varying task difficulties.
5.3 Comparison with other Frameworks.
@@
To further highlight the advantages of the AppAgentX framework, we compare its performance to other state-of-the-art frameworks on additional benchmarks, including Mobile-Bench Deng et al. (2024) and DroidTask Wen et al. (2023). As shown in Figure 2, AppAgentX outperforms AppAgent, in both task execution time and accuracy. Specifically, AppAgentX achieves significantly higher efficiency while maintaining higher task success rates across a broader range of applications.
Additionally, we compare the efficiency of AppAgentX against the Mobile-Agent-v2 framework Wang et al. (2024a) under different foundational LLMs, including GPT-4 OpenAI (2024), Claude 3.5 Sonnet Anthropic. (2024), and Gemini 1.5 Pro Gemini (2024). Table 3 shows that AppAgentX consistently outperforms Mobile-Agent-v2Wang et al. (2024a) (previous sota) in terms of efficiency, highlighting its ability to optimize performance across various LLMs.
@@
Figure 5: Qualitative Task Evaluation. This figure presents the qualitative results of path execution utilizing shortcuts on different applications (Gmail and Apple Music). It demonstrates the efficiency of AppAgentX in decreasing the utilization of LLM for per-step reasoning.
5.4 Qualitative Analysis
@@
In Figures 5, we show a qualitative analysis of some of the actions that used high-level actions to operate the screen. Despite the different user interfaces and actions performed by these applications, AppAgentX successfully completed the given task.
@@ 6 Conclusion
Limitations
As the GUI Agent framework continues to evolve, the recognition and processing of screen content remain a primary bottleneck in current research. While target detection and segmentation techniques in computer vision have reached a relatively mature stage, the detection of interactive elements on screens is still in its early development and lacks sufficient stability and generalization capability. This limitation can introduce biases in the Agent’s perception of the environment, ultimately affecting the accuracy of its interaction decisions.
In this study, we attempt to address this challenge by employing a dynamic matching mechanism in conjunction with the state-of-the-art screen element detection method. However, due to the diversity of screen elements, the complexity of layouts, and the variability across different application interfaces, our approach still has limitations and does not fundamentally overcome the bottleneck of screen content understanding and localization. Therefore, future research should focus on developing more robust screen content understanding techniques, such as integrating large-scale pre-trained visual models, joint image-text modeling, and other advanced methods, to enhance the adaptability and generalization capability of GUI Agents in complex environments.
Appendix A Appendix
@@ A.1 Chain Design Details
A.2 Execution Process Flowchart
Figure 6 presents a detailed flowchart that delineates the task processing workflow. The primary components of this workflow include page analysis, element comparison, matching, and memory storage.
@@ A.2 Execution Process Flowchart
A.3 Reliability Analysis of Shortcut Designs
In this study, we implement a fault-tolerant design for high-level actions within the evolutionary mechanism to enhance the overall robustness of the model. As depicted in Figure 6, the system utilizes a hierarchical action framework comprising high-level nodes and basic action spaces. During the operation, if the matching retrieval process of a high-level node fails, the system seamlessly transitions to the corresponding basic action spaces. This fallback strategy ensures that the model remains operational even when high-level decision-making encounters issues, thereby improving its resilience to errors and unexpected scenarios.

### v2

# AppAgentX: Evolving GUI Agents as Proficient Smartphone Users

@@ Abstract
Wenjia Jiang1,2, Yangyang Zhuang1, Chenxi Song1, Xu Yang3, Joey Tianyi Zhou4 Chi Zhang1, 1Westlake University, 2Henan University, 3Southeast University 4A*STAR {jiangwenjia, chizhang}@westlake.edu.cn, https://appagentx.github.io/
@@ 2 Related Works
LLMs as GUI Agents. Beyond general agents, LLMs have also enhanced GUI automation, surpassing traditional script-based methods in flexibility and adaptability. WebVoyagerHe et al. (2024), AppAgentZhang et al. (2023), and MobileAgentWang et al. (2024b) leverage multimodal perception for interactive interfaces, while UFOZhang et al. (2024a), AutoGLMLiu et al. (2024), and MMAC-CopilotSong et al. (2025) improve cross-platform adaptability and multi-agent collaboration. To refine UI understanding, OmniParserLu et al. (2024) and Ferret-UIYou et al. (2024a) enhance element recognition, while TinyClickPawlowski et al. (2024) and CogAgentHong et al. (2023) improve interaction precision and vision-based task execution. Additionally, WebGUMFuruta et al. (2024), MobileVLMSChu et al. (2024), and DigiRLBai et al. (2024) optimize performance in dynamic web and mobile environments. However, most existing agents rely on static training data and lack continuous adaptation. To address this limitation, we propose AppAgentX, integrating task trajectory memory to enhance efficiency and adaptability in GUI automation.
@@ 4.2 Evolutionary Mechanism
$\mathcal{A}_{\mathrm{evolved}}=\mathcal{A}_{\mathrm{basic}}\cup\{\tilde{a}\}.$ (2)
@@
4.3 Dynamic Action Execution
Following the evolution of the action space, the agent can now select high-level actions from the expanded action space to efficiently execute tasks.
4.3.1 Execution Process
The high-level execution process begins after the agent captures a screenshot of the current page. At this stage, the system then matches the parsed elements on the page to the stored element nodes in memory, by comparing their visual embeddings. Next, we check whether these identified element nodes are associated with any shortcut nodes. If an association between the element nodes and shortcut nodes exists, the system leverages the LLM to determine whether the corresponding high-level actions can be executed. This decision is made by inspecting the description of the shortcut node in conjunction with the current task context. If the conditions for executing the high-level action are met, the LLM generates an action execution template. This template includes the sequence of low-level actions to be executed by the shortcut node, along with the corresponding function arguments necessary for each action. In cases where multiple elements on the page are associated with shortcut nodes, the system prioritizes executing the actions based on the order of matching elements. This order is determined by their execution sequence, ensuring that the actions are performed in a logical and efficient manner. As the sequence actions progresses, partially repetitive operations or even the entire task can be completed more efficiently through the use of high-level actions.
4.3.2 Fallback Strategy
To ensure robustness and task completion reliability, we introduce a fallback strategy that allows the agent to dynamically recover from failures. If the conditions for executing a high-level action are not met, the agent will default to selecting actions from the basic action space $\mathcal{A}$. Additionally, in cases where execution errors occur due to incorrect shortcut node matching or unexpected UI responses, the agent reverts to using actions from the basic action space as a fallback. This ensures that the agent can still operate effectively, even when high-level abstractions cannot be applied to the current task.
@@ 5.1 Experimental Setup
Table 1: Analysis of Different Components in AppAgentX. This table compares the performance differences resulting from the different designs with the baseline. In that table for the GPT-4o approach, we use direct LLM invocation. Both our memory design and evolution mechanism can improve success rate and efficiency.
[…]
GPT-4o None Basic 10.8 26 6.72 16.9%
AppAgent Element Basic 9.3 24 8.46 69.7%
AppAgentX Chain Basic 9.1 23 9.26 70.8%
AppAgentX Chain Basic+Evolve 5.7 16 4.94 71.4%
@@
AndroidWorld Rawles et al. (2025): AndroidWorld is a fully functional and dynamic Android benchmark that supports 116 programmatic tasks across 20 widely used real-world Android applications.
A3 (Android Agent Arena) Chai et al. (2025): It consists of 201 tasks derived from 20 widely used third-party applications, covering common user scenarios. The benchmark provides a comprehensive evaluation mechanism, which significantly accelerates our experimental work. Implementation details. The implementation of our framework leverages several key platforms. For the foundational LLM, we selected GPT-4o OpenAI (2024) as the default model unless otherwise stated. The LangGraph framework LangChain (2024) is used as the agent platform, providing essential features for LLM input-output parsing and process control. To implement the memory mechanism, we integrated Neo4j Neo4j (2024) for graph-based storage and retrieval and Pinecone Pinecone (2024) for vector search. For feature matching, we employed cosine similarity with embeddings derived from ResNet-50 He et al. (2015), which enables effective task representation and retrieval. All experiments were conducted using the Android Debug Bridge (ADB), enabling on-device evaluations for mobile applications.
@@ 5.2 Experimental Analysis
Comparative Analysis. A comparison of our experimental results with model variants is shown in Table 1. We report the results on the AppAgent benchmark. We begin with a baseline model that does not incorporate any memory mechanism and progressively introduce our proposed enhancements. The first model variant integrates a module that records only information about elements, providing a basic form of memory. Subsequently, we introduce a more advanced memory design that maintains a structured memory chain, enabling the model to retain a more comprehensive representation of past interactions. Finally, we augment the system with an evolutionary mechanism that expands the action space to include high-level actions, further optimizing task execution.
@@
Table 2: Comparison with AppAgent on Large Benchmarks. This table evaluates the efficiency and accuracy of different frameworks on benchmarks containing a large number of tasks.
[…]
DroidTaskWen et al. (2024) 158 AppAgent 106.24 11.5 46.3%
AppAgentX 56.29 5.1 88.2%
AndroidWorldRawles et al. (2025) 116 AppAgent 147.17 18.9 41.7%
AppAgentX 59.74 6.2 62.5%
A3 (Android Agent Arena)Chai et al. (2025) 201 AppAgent 134.67 19.2 10.3%
AppAgentX 48.12 4.7 39.3%
@@
To further highlight the advantages of the AppAgentX framework, we compare its performance to other state-of-the-art frameworks on additional benchmarks, including DroidTask Wen et al. (2024), Android Agent ArenaChai et al. (2025) and AndroidWorldRawles et al. (2025). As shown in Figure 2, AppAgentX outperforms AppAgent, in both task execution time and accuracy. Specifically, AppAgentX achieves significantly higher efficiency while maintaining higher task success rates across a broader range of applications.
Figure 4: Comparison of Average Execution Time per Step. This figure presents the average execution time per steps across different LLMs and frameworks.
Additionally, we compare the execution efficiency of AppAgentX against two other frameworks Wang et al. (2024a); Zhang et al. (2023) across several prominent foundational LLMs, including GPT-4o OpenAI (2024), Claude 3.5 Sonnet Anthropic. (2024), and Gemini 1.5 Pro Gemini (2024). While acknowledging that certain discrepancies in the experimental setup may exist, we focus our evaluation primarily on execution time, as it serves as a practical and reproducible metric for assessing efficiency. As shown in Table 4, AppAgentX consistently demonstrates faster per-step completion times compared to the previous state-of-the-art Wang et al. (2024a) and AppAgent, across multiple LLM backends. These results suggest improved efficiency and better alignment with real-world deployment constraints.
Task difficulties analysis.
@@
Figure 6: Qualitative Results. This figure presents the path execution utilizing shortcuts on Gmail. It demonstrates the efficiency of AppAgentX in decreasing the utilization of LLM for per-step reasoning.
To investigate the performance of AppAgentX across tasks of varying complexity, we categorize the tasks based on their ground-truth operation steps, annotated through expert demonstrations. Tasks with up to 5 steps are considered short, tasks with 6 to 10 steps are classified as medium-length, and tasks requiring more than 10 steps are categorized as long tasks. As shown in Figure 5, the violin plot provides an intuitive visualization of how task complexity affects performance. Notably, our method demonstrates a clear advantage in terms of task execution time as the task complexity increases. To account for random fluctuations, we conducted one-tailed paired t-tests Ross and Willson (2017) on each data group to verify that AppAgentX significantly outperforms the baseline in efficiency, with all p-values falling below 0.05. Our model consistently outperforms the baseline across all task lengths, resulting in reduced task completion times. This finding reinforces the robustness and efficiency of our approach, confirming its stability under varying task difficulties.
Table 3: Ablation Study of Different Screen Perceptrons and Action Spaces. This table compares task success rates and times for different screen parsers and two action space variants: BAS (Basic Action Space) and FAS (Full Action Space).
[…]
Omniparser + BAS 209.3s 70.8%
Omniparser + FAS 91.2s 71.4%
Ferret-UI + BAS 201.7s 68.1%
Ferret-UI + FAS 92.5s 70.6%
Ablation Study. To further evaluate the performance gains of our agent, we substitute different screen perceptrons to assess the effectiveness of our evolutionary approach. Specifically, OmniParserLu et al. (2024) is employed as the default visual element extractor to locate UI components on the screen. In this study, we replace it with Ferret-UIYou et al. (2024b), a component with equivalent functionality, and compare the baseline action space with the evolved action space to examine their respective impacts.
As shown in Table 3, the Full Action Space significantly improves task performance compared to the Basic Action Space, significantly reducing execution time while simultaneously improving success rates. These results validate the effectiveness of the evolved action space.
@@
Qualitative Analysis. In Figures 6, we show a qualitative analysis of some of the actions that used high-level actions to operate the screen. On the left side we label the various processes in the relevant execution operations, from the retrieval of tasks to the matching of shortcut nodes. Despite the different user interfaces and actions performed by these applications, AppAgentX successfully completed the given task.
@@ 6 Conclusion
Appendix A Design and Execution Details
This section presents the detailed design of the proposed chain structure and the overall task execution workflow in our system.
@@ A.1 Chain Design Details
Together, these types of edges, along with the nodes described in the main text, constitute the structure of the complete memory store. This modeling approach aligns with the relationships between page jumps and containment, and offers a solid foundation for analyzing the evolution of actions.
A.2 Execution Process Details
@@
Figure 7 presents a detailed flowchart that delineates the task processing workflow. The primary components of this workflow include page analysis, element comparison, matching, and memory storage. As depicted in Figure 7, the system utilizes a hierarchical action framework comprising high-level nodes and basic action spaces. During the operation, if the matching retrieval process of a high-level node fails, the system seamlessly transitions to the corresponding basic action spaces.
Appendix B Detailed Numerical Results
This section provides the numerical analysis results of various shortcut designs, including evaluations of robustness and statistical tests to validate the effectiveness of the proposed system design.
To supplement the grouped bar chart presented in Figure 4, this subsection provides the exact per-step execution time (in seconds) for each evaluated framework (MobileAgent2, AppAgent, and AppAgentX) across different LLM backends. These numerical results are intended to enhance reproducibility, facilitate further analysis, and offer a precise basis for comparing execution efficiency across systems.
LLM Previous SOTA AppAgent Ours
Gemini-1.5-Pro 25.6 15.2 12.4
Claude-3.5 38.4 13.8 11.4
GPT-4o 43.5 20.3 17.5
Table 4 lists the detailed values corresponding to each bar in the figure.
Table 5: Statistical Test Results for Task Completion Times. Results from one-tailed paired t-tests.
Task Length Mean Difference (s) t-value p-value
Short Task -12.4 -3.45 <0.01
Medium Task -30.7 -5.12 <0.001
Long Task -75.3 -6.89 <0.001
Figure 8: Qualitative Results. This figure presents the path execution utilizing shortcuts on Apple Music. It demonstrates the efficiency of AppAgentX in decreasing the utilization of LLM for per-step reasoning.
To reinforce the observed differences in task completion times, we complemented the analysis with detailed results from one-tailed t-tests, as summarized in Table 5. These results address data not fully elaborated upon in the main text. The negative t-values reported in this table indicate the presence of a directional effect, where AppAgentX consistently outperforms the baseline in reducing task completion times. Specifically, the negative sign reflects that the mean completion time for AppAgentX is significantly lower than that of the baseline across all task lengths. This directional result aligns with our hypothesis and further validates the consistent efficiency advantage of AppAgentX.
Appendix C Additional Quantitative Analysis
To further illustrate the behavior of our system in real-world settings, we include a qualitative example in Figure 8. This example showcases the execution path for a task in Apple Music, highlighting how AppAgentX leverages existing shortcuts to complete the task more efficiently. By using pre-defined shortcuts, the system reduces the reliance on large language model (LLM) reasoning at each step, thereby improving both speed and resource usage.